OpenStudy (anonymous):

Solve system using elimination. 6x+3y=-12 6x+2y=-4

OpenStudy (anonymous):

6x+3y=-12 -6x-2y=4 ---> y=-8 and x=???

OpenStudy (anonymous):

I'm sorry, I don't understand how you got y=-8.

OpenStudy (anonymous):

\[6x + 3y = -12 \rightarrow 6x = -12 - 3y\]\[6x + 2y = -4 \rightarrow 6x = -4 -2y\]now we have 2 equation that one side of them is 6x, so other sides are equal!\[-12-3y = -4 - 2y \rightarrow -3y + 2y = -4 + 12 \rightarrow -y = 8\]so \[y = -8\] can you continue and solving system?

OpenStudy (anonymous):

So if that's how I find the y, do I do the opposite for the X? Like separate the y from the x and normal number?

OpenStudy (anonymous):

you should plug in -8 for y and then find x

OpenStudy (anonymous):

So 6x-16=-4?

OpenStudy (anonymous):

ohom, and 6x = 12 and x =2 ;)

OpenStudy (anonymous):

Okay, thank you so much.

OpenStudy (anonymous):

yw ;)

OpenStudy (anonymous):

\[y=-8 and x=2\]